# C2:Cross learning module enhanced decision transformer with Constraint-aware loss for auto-bidding

### v1

@@ Abstract.
Decision Transformer (DT) shows promise for generative auto-bidding by capturing temporal dependencies, but suffers from two critical limitations: insufficient cross-correlation modeling among state, action, and return-to-go (RTG) sequences, and indiscriminate learning of optimal/suboptimal behaviors. To address these, we propose C2, a novel framework enhancing DT with two core innovations: (1) a Cross Learning Block (CLB) via cross-attention to strengthen inter-sequence correlation modeling; (2) a Constraint-aware Loss (CL) incorporating budget and Cost-Per-Acquisition (CPA) constraints for selective learning of optimal trajectories. Extensive offline evaluations on the AuctionNet dataset demonstrate consistent performance gains (up to 3.23% over state-of-the-art GAVE) across diverse budget settings; ablation studies verify the complementary synergy of CLB and CL, confirming C2’s superiority in auto-bidding. The code for reproducing our results is available at: https://github.com/Dingjinren/C2.
@@
journalyear: 2026copyright: acmlicensedconference: arXiv Preprint (Unsubmitted to any conference); 2026; arXiv.orgbooktitle: arXiv Preprint (Unsubmitted to any conference) (arXiv Preprint), 2026, arXiv.orgdoi: 10.48550/arXiv.26XX.XXXXXXisbn: 978-1-4503-XXXX-X/26/01ccs: Information systems Computational advertising
@@ 4.2. Implementation Details
All experiments are executed on NVIDIA H800 GPUs with a fixed batch size of 128 and a maximum of 10,000 training iterations to ensure sufficient convergence; model parameters are optimized using the AdamW optimizer (Loshchilov, 2017) with a learning rate of $10^{-5}$, which is empirically determined to balance training stability and convergence speed, and specifically on the AuctionNet dataset, we set the hyperparameter $\lambda$ in Eq. (20) to 10 to appropriately weight the contribution of RTG Loss within the total loss function.

### v2

@@ Abstract.
Decision Transformer (DT) shows promise for generative auto-bidding by capturing temporal dependencies, but suffers from two critical limitations: insufficient cross-correlation modeling among state, action, and return-to-go (RTG) sequences, and indiscriminate learning of optimal/suboptimal behaviors. To address these, we propose C2, a novel framework enhancing DT with two core innovations: (1) a Cross Learning Block (CLB) via cross-attention to strengthen inter-sequence correlation modeling; (2) a Constraint-aware Loss (CL) incorporating budget and Cost-Per-Acquisition (CPA) constraints for selective learning of optimal trajectories. Extensive offline evaluations on the AuctionNet dataset demonstrate consistent performance gains (up to 3.23% over state-of-the-art method) across diverse budget settings; ablation studies verify the complementary synergy of CLB and CL, confirming C2’s superiority in auto-bidding. The code for reproducing our results is available at: https://github.com/Dingjinren/C2.
@@
journalyear: 2026copyright: acmlicensedconference: arXiv Preprint; 2026; arXiv.orgbooktitle: arXiv Preprint (arXiv Preprint), 2026, arXiv.orgdoi: 10.48550/arXiv.2601.20257isbn: 978-1-4503-XXXX-X/26/01ccs: Information systems Computational advertising
@@ 4.2. Implementation Details
All experiments are executed with a fixed batch size of 128 and a maximum of 10,000 training iterations to ensure sufficient convergence; model parameters are optimized using the AdamW optimizer (Loshchilov, 2017) with a learning rate of $10^{-5}$, which is empirically determined to balance training stability and convergence speed, and specifically on the AuctionNet dataset, we set the hyperparameter $\lambda$ in Eq. (20) to 10 to appropriately weight the contribution of RTG Loss within the total loss function.